# Positive intermediate Ricci curvature on cohomogeneity one manifolds in low dimensions

### v1

@@ Abstract.
We explore existence of invariant metrics with positive intermediate Ricci curvature on closed, low-dimensional cohomogeneity one manifolds. For a certain cohomogeneity one $\mathsf{Spin}(4)$-action on $S^{3}\times\mathbb{C}\mathrm{P}^{2}$, we construct an invariant metric with positive $4$-intermediate Ricci curvature and show it cannot admit one with positive $3$-intermediate Ricci curvature. We establish similar obstructions for several other cohomogeneity one manifolds.
@@ 1. Introduction
The study of Riemannian manifolds with positive sectional curvature has been a central problem in Riemannian geometry since the creation of the subject. While there are several topological obstructions to such manifolds, the list of known examples is very limited, with the rank one symmetric spaces remaining the only known simply connected examples in dimensions larger than $24$. An area of focus receiving increased attention lately has been to investigate positive $k^{th}$-intermediate Ricci curvature ($\operatorname{Ric}_{k}>0$), a condition that interpolates between positive sectional curvature ($k=1$) and positive Ricci curvature ($k=n-1$) on an $n$-manifold (see Definition 2.1). Motivating goals in this area include extending results known for positively curved manifolds (e.g [She93, Wil97, GW18, GW20, GW22, Mou22, KM24]), and constructing new examples which do not admit positive sectional curvature (e.g[AQZ20, DGM23, DDGR24, RW23a, RW22, RW23b, RW23c]).
Focusing on closed, simply connected cohomogeneity one manifolds, we have a complete classification of those with positive sectional curvature except in dimension seven, in which there are two families of possible candidates [Ver02, Ver04, GWZ08, VZ14]. (For an extension of this classification to quasipositive curvature, see [Wul24].) On the other hand, any cohomogeneity one manifold with finite fundamental group admits an invariant metric with positive Ricci curvature (see [GZ02]). This naturally leads to the question: What is the smallest $k$ for which a given cohomogeneity one manifold admits an invariant metric with $\operatorname{Ric}_{k}>0$?
@@
For a description of this action, see Equation (3.1) below. Taking the quotient of $S^{3}\times\mathbb{C}\mathrm{P}^{2}$ with the metric from Theorem 1.1 by the Hopf action, the Gray-O’Neill formula (Lemma 2.3) implies the following:
Corollary 1.2.
@@
In [VZ14], Verdiani and Ziller applied an argument involving transverse families of Jacobi fields to rule out existence of an invariant metric with $\operatorname{Ric}_{2}>0$ on a family of $7$-dimensional cohomogeneity one manifolds; see Example 4.1. We follow similar techniques to prove the last statement in Theorem 1.1. Taking this method further using the classification of low-dimensional cohomogeneity one manifolds in [Hoe10], we also prove the following:
Theorem 1.3.
@@
    $\{(1,1)\}\subset\{(e^{i\theta},1),(e^{jd\theta},e^{i\theta})\}\subset S^{3}\times S^{1}.$
@@
This paper is organized as follows: In Section 2, we collect tools and observations used throughout the paper. Section 3 is devoted to the construction of the metric with $\operatorname{Ric}_{4}>0$ on $S^{3}\times\mathbb{C}\mathrm{P}^{2}$ described in Theorem 1.1, while the proof of Theorem 1.3, as well as that of the second claim in Theorem 1.1, is given in Section 4. Throughout the paper, we follow the notational conventions established in [Hoe10].
@@ 2.1. Intermediate Ricci curvature
This metric is a Cheeger deformation of the product action via the action of the subgroup $S^{3}\times\{1\}\times\{1\}$. (See Section 2.4 for the definition of Cheeger deformation and Example 3.1 for why this metric has $\operatorname{Ric}_{2}>0$.) Letting $S^{1}$ denote the complex circle $\{e^{i\theta}\}$ in $S^{3}$, the quotient of $S^{3}\times S^{3}$ by the free action of the subgroup $\{1\}\times\{1\}\times S^{1}$ is diffeomorphic to $S^{3}\times S^{2}$, and it follows from Lemma 2.3 that this space inherits a metric with $\operatorname{Ric}_{2}>0$. This metric is invariant under the induced transitive action by $S^{3}\times S^{3}\times\{1\}$. On the other hand, the quotient of $S^{3}\times S^{3}$ by the subgroup $\{1\}\times S^{1}\times S^{1}$ provides a metric with $\operatorname{Ric}_{2}>0$ on $S^{2}\times S^{2}$. This metric is invariant under the induced cohomogeneity one action by $S^{3}\times\{1\}\times\{1\}$, which is equivalent to the diagonal action of $\mathsf{SO}(3)$ on $S^{2}\times S^{2}$.
@@ 2.2. Cohomogeneity one manifolds
Let $M$ be a cohomogeneity one manifold with the group diagram $H\subset\{K_{-},K_{+}\}\subset G$. A cohomogeneity one $G$-manifold $N$ is equivariantly diffeomorphic to $M$ if and only if its group diagram is of the form (up to switching $K_{-}$ and $K_{+}$):
@@ 2.3. Curvature restrictions on Jacobi fields
Suppose $(M^{n},g)$ is a compact, cohomogeneity one $G$-manifold with $M/G=[0,L]$. Let $\gamma:[0,L]\to M$ be a unit speed horizontal geodesic that begins at a singular orbit $N=G(\gamma(0))$. Consider the $(n-1)$-dimensional vector space $\Lambda$ of Jacobi fields along $\gamma$ generated by the $G$-action. For all $t\in(0,L)$, we have $\operatorname{span}\{\gamma^{\prime}(t)\}^{\perp}=\{J(t):J\in\Lambda\}$, and the Riccati operator $S_{t}:\operatorname{span}\{\gamma^{\prime}(t)\}^{\perp}\to\operatorname{span}\{\gamma^{\prime}(t)\}^{\perp}$ is defined by $S_{t}(J(t))=J^{\prime}(t)$. For $t=0$ or $L$, we have $\operatorname{span}\{\gamma^{\prime}(t)\}^{\perp}=\{J(t):J\in\Lambda\}\oplus\{J^{\prime}(t):J\in\Lambda,J(t)=0\}$ (see [GW09, Section 1.7]). Moreover, by Lemmma 1.3 in [GW22], the equation $S_{t}(J(t))=J^{\prime}(t)$ provides a well-defined Riccati operator for these values of $t$ as well. Since $\Lambda$ consists of variational fields for the variation of $\gamma$ by geodesics that leave $N$ orthogonally at $t=0$, it follows that the Riccati operator $S_{t}$ is self-adjoint for all $t$ (see Lemma 1.4 in [GW22]). $\Lambda$ is an example of a Lagrangian family of normal Jacobi fields. In other words, with respect to the symplectic form $\omega(J_{1},J_{2})=g(J_{1}^{\prime},J_{2})-g(J_{1},J_{2}^{\prime})$ defined on the $(2n-2)$-dimensional space of all Jacobi fields orthogonal to $\gamma$, $\Lambda$ is equal to its own symplectic complement.
@@
{±$\{\pm 1\}\subset\{e^{i\theta}\},\{e^{i\theta}\}\subset S^{3},$
{±$\{\pm 1\}\subset\{e^{i\theta}\},\{e^{j\theta}\}\subset S^{3}.$
@@ 2.4. Cheeger deformation
In [Che73], Cheeger developed a method for deforming Riemannian metrics in the presence of isometric group actions, now referred to as a Cheeger deformation. In this section, we collect some general facts about Cheeger deformations that will be used in Section 3 to construct the metric in Theorem 1.1.
[…]
@@ 3. Constructions
The main goal of this section is to prove that $S^{3}\times\mathbb{C}\mathrm{P}^{2}$ admits a metric with $\operatorname{Ric}_{4}>0$ as in Theorem 1.1. We begin with a description of the metric with $\operatorname{Ric}_{2}>0$ on $S^{3}\times S^{3}$ referenced in Example 2.4. Wilking described the zero curvature planes for this metric in [Wil02, Section 5], from which it can be inferred that this metric has $\operatorname{Ric}_{2}>0$. The second author with Domínguez-Vázquez and González-Álvaro extended this construction to products of the form $G\times G$, where $G$ is a semisimple Lie group (see [DGM23, Theorem E]). We describe the construction here to provide a simple example of how the tools from Section 2.4 are used.
@@
Consider $S^{3}$ with the standard round metric $g_{S^{3}}$, and equip $S^{3}\times S^{3}$ with the product metric $g\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=g_{S^{3}}\oplus g_{S^{3}}$. Let $g_{\ell}$ denote the Cheeger deformation of $g$ with respect to the $S^{3}$-action given by
@@
Now, we will construct the metric referenced in Theorem 1.1. Let $\phi:S^{3}\to\mathsf{SO}(3)$ denote the standard covering map, and think of $\mathsf{SO}(3)$ as a subgroup of $\mathsf{SU}(3)$. Then $S^{3}\times\mathbb{C}\mathrm{P}^{2}$ admits a cohomogeneity one action of $S^{3}\times S^{3}$ given by
@@
This cohomogeneity one manifold belongs to the the family $P_{A}^{7}$ described in Proposition 4.6 with $p_{-}=p_{+}=q_{-}=q_{+}=1$ (see [Hoe10, p.174]), where we show that these manifolds cannot admit an invariant metric with $\operatorname{Ric}_{3}>0$. Hence the following result is optimal:
Proposition 3.2.
$S^{3}\times\mathbb{C}\mathrm{P}^{2}$ admits a metric with $\operatorname{Ric}_{4}>0$ that is invariant under the cohomogeneity one $(S^{3}\times S^{3})$-action given by (3.1).
Proof.
We start by considering the lift of the $(S^{3}\times S^{3})$-action from (3.1) to $S^{3}\times S^{5}\subset\mathbb{H}\oplus\mathbb{C}^{3}$, which is given by
$(a,b)*(p,z)=(apb^{-1},\phi(b)z).$
@@
Consider $S^{3}\times S^{5}$ equipped with the standard product metric $g=g_{S^{3}}\oplus g_{S^{5}}$. Then the $(S^{3}\times S^{3})$-action is by isometries of $g$. Moreover, the metric $g_{\ell}$ obtained by the Cheeger deformation of $g$ via the $S^{3}$-action is invariant under the Hopf action. Therefore, there exists a Riemannian metric $\check{g}_{\ell}$ on $S^{3}\times\mathbb{C}\mathrm{P}^{2}$ such that the quotient map $(S^{3}\times S^{5},g_{\ell})\to(S^{3}\times\mathbb{C}\mathrm{P}^{2},\check{g}_{\ell})$ is a Riemannian submersion. We claim that $\check{g}_{\ell}$ has $\operatorname{Ric}_{4}>0$.
@@ 4. Obstructions
In this section, we investigate cohomogeneity one manifolds in dimensions 55557777, and show that some of them cannot admit invariant metrics with $\operatorname{Ric}_{k}>0$ for certian values of $k$. We are motivated by the following example:

### v2

@@ Abstract.
We explore existence of invariant metrics with positive intermediate Ricci curvature on closed, low-dimensional cohomogeneity one manifolds. For a certain cohomogeneity one $\mathsf{Spin}(4)$-action on $S^{3}\times\mathbb{C}\mathrm{P}^{2}$, we construct an invariant metric with positive $4^{\mathrm{th}}$-intermediate Ricci curvature and show it cannot admit one with positive $3^{\mathrm{rd}}$-intermediate Ricci curvature. We further establish similar symmetry obstructions to positive curvature for $S^{3}\times S^{3}$, $S^{3}\times S^{4}$, and several families of cohomogeneity one manifolds.
@@ 1. Introduction
The study of Riemannian manifolds with positive sectional curvature has been a central problem in Riemannian geometry since the creation of the subject. While there are several topological obstructions to such manifolds, the list of known examples is very limited, with the rank one symmetric spaces remaining the only known simply connected examples in dimensions larger than $24$. An area of focus receiving increased attention lately has been the investigation positive $k^{th}$-intermediate Ricci curvature ($\operatorname{Ric}_{k}>0$), a condition that interpolates between positive sectional curvature ($k=1$) and positive Ricci curvature (k=n1k=n-1) on an $n$-manifold (see Definition 2.1). Motivating goals in this area include extending results known for positively curved manifolds (e.g [She93, Wil97, GW18, GW20, GW22, Mou22, KM24]), and constructing new examples which do not admit positive sectional curvature (e.g[AQZ20, DGM23, DDGR24, RW23a, RW22, RW23b, RW23c]).
Focusing on closed, simply connected cohomogeneity one manifolds, we have a complete classification of those with positive sectional curvature except in dimension seven, in which there are two families of possible candidates [Ver02, Ver04, GWZ08, VZ14]. (For an extension of this classification to quasipositive curvature, see [Wul24].) On the other hand, any cohomogeneity one manifold with finite fundamental group admits an invariant metric with positive Ricci curvature [GZ02]. This naturally leads to the question: What is the smallest $k$ for which a given cohomogeneity one manifold admits an invariant metric with $\operatorname{Ric}_{k}>0$?
@@
For a description of this action, see Equation (3.1) below. The metric we construct is a Cheeger deformation by the action of an $S^{3}$ subgroup; for details see the exposition preceding Proposition 3.2. The Hopf action on the $S^{3}$-factor is by isometries of this metric, thus the Gray-O’Neill formula (Lemma 2.3) applied to the quotient map $S^{3}\times\mathbb{C}\mathrm{P}^{2}\to S^{2}\times\mathbb{C}\mathrm{P}^{2}$ gives the following:
Corollary B.
@@
Since our metrics in Theorem A and Corollary B are of cohomogeneity one and two, respectively, it is natural to ask the following:
Question 1.1.
Does $S^{3}\times\mathbb{C}\mathrm{P}^{2}$ or $S^{2}\times\mathbb{C}\mathrm{P}^{2}$ admit a homogeneous metric with $\operatorname{Ric}_{4}>0$?
In [VZ14], Verdiani and Ziller applied an argument involving transverse families of Jacobi fields to rule out existence of an invariant metric with $\operatorname{Ric}_{2}>0$ on a family of $7$-dimensional cohomogeneity one manifolds; see Example 4.1. We follow similar techniques to prove the last statement in Theorem A. Taking this method further using the classification of low-dimensional cohomogeneity one manifolds in [Hoe10], we also prove the following:
Theorem C.
@@
    $\{(1,1)\}\subset\{(e^{i\theta},1),(e^{jd\theta},e^{i\theta})\}\subset S^{3}\times S^{1},$
    where $d$ is even.
  2. (2)
@@
Notice that in Theorem C, we rule out $(S^{3}\times S^{1})$-invariant metrics with $\operatorname{Ric}_{2}>0$ on the $5$-manifolds $M_{d}^{5}$ and $Q_{A}^{5}$. According to [Hoe10, Theorem C], the only compact, simply connected cohomogeneity one manifolds in dimension five are diffeomorphic to the sphere $S^{5}$, the Wu manifold $\mathsf{SU}(3)/\mathsf{SO}(3)$, or either of the two $S^{3}$-bundles over $S^{2}$. Furthermore, each of these manifolds admits a cohomogeneity one action by $S^{3}\times S^{1}$. The round metric on $S^{5}$ has positive sectional curvature and is invariant under such an action. Also, $S^{3}\times S^{2}$ is known to admit an metric with $\operatorname{Ric}_{2}>0$ that is invariant under a cohomogeneity one $(S^{3}\times S^{1})$-action; see Example 2.4 below. Since it is currently unknown whether the Wu manifold or the non-trivial $S^{3}$-bundle over $S^{2}$ admit metrics with $\operatorname{Ric}_{2}>0$, we pose the following:
Question 1.2.
Does $\mathsf{SU}(3)/\mathsf{SO}(3)$ or the non-trivial $S^{3}$-bundle over $S^{2}$ admit a metric with $\operatorname{Ric}_{2}>0$ that is invariant under a cohomogeneity one $(S^{3}\times S^{1})$-action?
In dimension four, the only closed, simply connected cohomogeneity one manifolds are diffeomorphic to $S^{4}$, $\mathbb{C}\mathrm{P}^{2}$, $S^{2}\times S^{2}$, and $\mathbb{C}\mathrm{P}^{2}\#\overline{\mathbb{C}\mathrm{P}^{2}}$. While the first two admit metrics with positive sectional curvature and $S^{2}\times S^{2}$ admits a cohomogeneity one metric with $\operatorname{Ric}_{2}>0$ (see Example 2.4), the following is still open:
Question 1.3.
Does $\mathbb{C}\mathrm{P}^{2}\#\overline{\mathbb{C}\mathrm{P}^{2}}$ admit a metric with $\operatorname{Ric}_{2}>0$, perhaps invariant under a cohomogeneity one action?
This paper is organized as follows: In Section 2, we collect tools and observations used throughout the paper. Section 3 is devoted to the construction of the metric with $\operatorname{Ric}_{4}>0$ on $S^{3}\times\mathbb{C}\mathrm{P}^{2}$ described in Theorem A. The proof of Theorem C, as well as that of the second claim in Theorem A, is given in Section 4. Specifically, the proof of Theorem A is given in Propositions 3.2 and 4.6. Throughout the paper, we follow the notational conventions established in [Hoe10].
Acknowledgments
This work was initiated when the second author visited Clark University; he thanks the Department of Mathematics for their hospitality. The authors would like to thank Lee Kennard and Jason DeVito for their comments on a preliminary version of this article. The first author was supported by the AMS-Simons Travel Grant, while the second author was supported by NSF Award DMS-2202826 and the AMS-Simons Research Enhancement Grant for PUI Faculty.
@@ 2.1. Intermediate Ricci curvature
This metric is a Cheeger deformation of the product action via the action of the subgroup $S^{3}\times\{1\}\times\{1\}$. (See Section 2.4 for the definition of Cheeger deformation and Example 3.1 for why this metric has $\operatorname{Ric}_{2}>0$.) Letting $S^{1}$ denote the complex circle $\{e^{i\theta}\}$ in $S^{3}$, the quotient of $S^{3}\times S^{3}$ by the free action of the subgroup $\{1\}\times\{1\}\times S^{1}$ is diffeomorphic to $S^{3}\times S^{2}$, and it follows from Lemma 2.3 that this space inherits a metric with $\operatorname{Ric}_{2}>0$. This metric is invariant under the induced transitive action by $S^{3}\times S^{3}\times\{1\}$, and hence it is invariant under the cohomogeneity one action by $S^{3}\times S^{1}\times\{1\}$. On the other hand, the quotient of $S^{3}\times S^{3}$ by the subgroup $\{1\}\times S^{1}\times S^{1}$ provides a metric with $\operatorname{Ric}_{2}>0$ on $S^{2}\times S^{2}$. This metric is invariant under the induced cohomogeneity one action by $S^{3}\times\{1\}\times\{1\}$, which is equivalent to the diagonal action of $\mathsf{SO}(3)$ on $S^{2}\times S^{2}$.
@@ 2.2. Cohomogeneity one manifolds
Let $M$ be a cohomogeneity one manifold with the group diagram $H\subset\{K_{-},K_{+}\}\subset G$. A cohomogeneity one $G$-manifold NN is equivariantly diffeomorphic to $M$ if and only if its group diagram is of the form (up to switching $K_{-}$ and $K_{+}$):
@@ 2.3. Curvature restrictions on Jacobi fields
Suppose $(M^{n},g)$ is a compact, cohomogeneity one $G$-manifold with $M/G=[0,L]$. Let $\gamma:[0,L]\to M$ be a unit speed horizontal geodesic that begins at a singular orbit $N=G(\gamma(0))$. Consider the (n1)(n-1)-dimensional vector space $\Lambda$ of Jacobi fields along $\gamma$ generated by the $G$-action. For all $t\in(0,L)$, we have $\operatorname{span}\{\gamma^{\prime}(t)\}^{\perp}=\{J(t):J\in\Lambda\}$, and the Riccati operator $S_{t}:\operatorname{span}\{\gamma^{\prime}(t)\}^{\perp}\to\operatorname{span}\{\gamma^{\prime}(t)\}^{\perp}$ is defined by $S_{t}(J(t))=J^{\prime}(t)$. For $t=0$ or $L$, we have $\operatorname{span}\{\gamma^{\prime}(t)\}^{\perp}=\{J(t):J\in\Lambda\}\oplus\{J^{\prime}(t):J\in\Lambda,J(t)=0\}$ (see [GW09, Section 1.7]). Moreover, by Lemmma 1.3 in [GW22], the equation $S_{t}(J(t))=J^{\prime}(t)$ provides a well-defined Riccati operator for these values of $t$ as well. Since $\Lambda$ consists of variational fields for the variation of $\gamma$ by geodesics that leave $N$ orthogonally at $t=0$, it follows that the Riccati operator $S_{t}$ is self-adjoint for all $t$ (see Lemma 1.4 in [GW22]). $\Lambda$ is an example of a Lagrangian family of normal Jacobi fields. In other words, with respect to the symplectic form $\omega(J_{1},J_{2})=g(J_{1}^{\prime},J_{2})-g(J_{1},J_{2}^{\prime})$ defined on the (2n2)(2n-2)-dimensional space of all Jacobi fields orthogonal to $\gamma$, $\Lambda$ is equal to its own symplectic complement.
@@
$\{\pm 1\}\subset\{e^{i\theta}\},\{e^{i\theta}\}\subset S^{3},$
$\{\pm 1\}\subset\{e^{i\theta}\},\{e^{j\theta}\}\subset S^{3}.$
@@ 2.4. Cheeger deformation
In [Che73], Cheeger developed a method for deforming Riemannian metrics in the presence of isometric group actions, now referred to as a Cheeger deformation. In this section, we collect some general facts about Cheeger deformations that will be used in Section 3 to construct the metric in Theorem A.
Let $G$ be a compact group of isometries of $(M,g)$, and let $Q$ be a bi-invariant metric on $G$. Consider the one-parameter family of metrics $\{\bar{g}_{\ell}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\ell^{2}Q+g\}_{\ell>0}$ on $G\times M$. Note that $G$ acts freely and by isometries on $(G\times M,\bar{g}_{\ell})$ via
@@ 3. Constructions
The main goal of this section is to prove that $S^{3}\times\mathbb{C}\mathrm{P}^{2}$ admits a metric with $\operatorname{Ric}_{4}>0$ as in Theorem A. We begin with a description of the metric with $\operatorname{Ric}_{2}>0$ on $S^{3}\times S^{3}$ referenced in Example 2.4. Wilking described the zero curvature planes for this metric in [Wil02, Section 5], from which it can be inferred that this metric has $\operatorname{Ric}_{2}>0$. The second author with Domínguez-Vázquez and González-Álvaro extended this construction to products of the form $G\times G$, where $G$ is a semisimple Lie group (see [DGM23, Theorem E]). We describe the construction here to provide a simple example of how the tools from Section 2.4 are used.
@@
[…]
@@
Now, we will construct the metric referenced in Theorem A. Let $\phi:S^{3}\to\mathsf{SO}(3)$ denote the standard covering map, and think of $\mathsf{SO}(3)$ as a subgroup of $\mathsf{SU}(3)$. Then $S^{3}\times\mathbb{C}\mathrm{P}^{2}$ admits a cohomogeneity one action of $S^{3}\times S^{3}$ given by
@@
Consider the lift of this action to $S^{3}\times S^{5}\subset\mathbb{H}\oplus\mathbb{C}^{3}$, which is given by
(3.2) $(a,b)*(p,z)=(apb^{-1},\phi(b)z).$
@@
Consider $S^{3}\times S^{5}$ equipped with the standard product metric $g=g_{S^{3}}\oplus g_{S^{5}}$. Then the $(S^{3}\times S^{3})$-action from (3.2) is by isometries of $g$. Let $g_{\ell}$ denote the Cheeger deformation of $g$ via the $S^{3}$-action, as defined in Section 2.4. Note that because the Hopf action commutes with the $S^{3}$-action, $g_{\ell}$ is invariant under the Hopf action by Lemma 2.9. Therefore, there exists a Riemannian metric $\check{g}_{\ell}$ on $S^{3}\times\mathbb{C}\mathrm{P}^{2}$ such that the quotient map
(3.3) $(S^{3}\times S^{5},g_{\ell})\to(S^{3}\times\mathbb{C}\mathrm{P}^{2},\check{g}_{\ell})$
is a Riemannian submersion.
The space $S^{3}\times\mathbb{C}\mathrm{P}^{2}$ with the action in (3.1) belongs to the the family $P_{A}^{7}$ described in Proposition 4.6 with $p_{-}=p_{+}=q_{-}=q_{+}=1$ (see [Hoe10, p.174]), where we show that these manifolds cannot admit an invariant metric with $\operatorname{Ric}_{3}>0$. Hence the following result is optimal:
Proposition 3.2.
The metric $\check{g}_{\ell}$ from (3.3) on $S^{3}\times\mathbb{C}\mathrm{P}^{2}$ has $\operatorname{Ric}_{4}>0$, and it is invariant under the cohomogeneity one $(S^{3}\times S^{3})$-action given by (3.1).
Proof.
@@ 4. Obstructions
In this section, we investigate cohomogeneity one manifolds in dimensions 5577, and show that some of them cannot admit invariant metrics with $\operatorname{Ric}_{k}>0$ for certian values of $k$. We are motivated by the following example: